OpenStudy (ja1):

Find the value in simplest form:

OpenStudy (ja1):

\[\frac{ \sqrt24 }{ 4\sqrt{6} }\]

OpenStudy (ja1):

@hicstps what I need is help with this not an answer.

OpenStudy (ja1):

Besides that was wrong anyway

OpenStudy (unklerhaukus):

is that the sqrt of twenty four in the nuemerator, or sqrt two times four?

OpenStudy (ja1):

√24 not 2 * 4

OpenStudy (unklerhaukus):

\[\sqrt{24}=\sqrt6\times\sqrt 4\]

OpenStudy (unklerhaukus):

the sqrt 6 will cancel

OpenStudy (ja1):

Ok yes i have that so far

OpenStudy (ja1):

Ohhhh right so I have this: \[\frac{ 2\sqrt{6} }{ 4\sqrt{6} }\]

OpenStudy (ja1):

So sqrt{6} will cancel

OpenStudy (unklerhaukus):

yes, yes

OpenStudy (ja1):

So I would then have 2/4 which makes .50?

OpenStudy (unklerhaukus):

that's right!\[\color{red}\checkmark\]

OpenStudy (ja1):

Awesome thanks I had forgotten that part of it :D

OpenStudy (ja1):

Well actually the answer is 1/2 but still right xD

OpenStudy (unklerhaukus):

yeah its a half